OpenStudy (anonymous):

log2(12b-21)-log2(b2-3)=2

OpenStudy (anonymous):

is it base "2"? or a number "2"?

OpenStudy (caozeyuan):

\[\log_{2} (12b-21) -\log_{2}(b ^{2}-3) = 2\]

OpenStudy (caozeyuan):

\[\log_{2} \frac{ 12b-21 }{ b ^{2}-3 }=2\]

OpenStudy (anonymous):

Apply the "Laws of Logarithm" and express in Exponential Form to find the value of "b". :) continuation of "caozeyuan's solution": exponential form \[\frac{ 12b-21 }{ b ^{2}-3 }=2^{2}\] cross-multiply \[12b-21=4b ^{2}-12\] combine like terms and equate to zero \[4b ^{2}-12b+9=0\] look for factors of the trinomial. :)

OpenStudy (anonymous):

It's a perfect square trinomial: \[(2b-3)^{2}\] zero property \[2b-3=0\]\[2b=3\]\[b=\frac{ 3 }{ 2 }\]that's the answer. :)